Given the equation, 9x + 3y =-6, write a second linear equation to create a system that has exactly only one solution. How do you do this ?
Answer by richwmiller(17219)

Lines that intersect have only one solution.
Lines that are perpendicular intersect.
So one way is to write an equation that is perpendicular.
Lines that are perpendicular have slopes that multiply to -1
9x + 3y =-6
is the same line as
y = -3 x-2 with slope
-3
1/3 would be the slope of a line perpendicular.
y = x/3-2 or y = 1/3*x-2